THERMODYNAMIC QUANTITIES OF CARBON NANOTUBES FROM SPECTROSCOPIC DATA

K.N. UTTAM, RENU SINGH and \underline{PAVITRA TANDON}, Saha’s Spectroscopy Laboratory, Department of Physics, University of Allahabad, Allahabad (U.P.), India.

Thermodynamical quantities like Free Energy, Enthalpy, Entropy and Specific heat of Carbon nanotubes have been calculated from spectroscopic data using partition theory. The results have been calculated in the temperature range 50-2150 K. The plot of the variation of the heat capacity with temperature have been studied and obtained results have been discussed in the light of spectroscopic data.